Circular No. 7413

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cc IN CGCG 140-014 Y. Sato and W. D. Li, University of California at Berkeley, on behalf of

the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Apr. 30.5 (limiting mag ∼ 19.0) and May 1.5 UT (limiting mag ∼ 20.5), both showing the new star at mag ∼ 18.5. SN 2000cc was also present on an unfiltered image taken on Apr. 26.5 at about the same brightness, and the new object is located at α = 17^{h}23^{m}14^{s}.34, δ = +23^{o}40′18′′.7 (equinox 2000.0), which is 5′′.9 east and 9′′.1 south of the nucleus of CGCG 140-014. Unfiltered images of the field taken on 2000 Apr. 10.5 (limiting mag ∼ 19.0) and 1998 July 16.3 (limiting mag ∼ 19.5) showed nothing at the position of the new object.

SUPERNOVA 2000ca IN ESO 383-32 G. Aldering and A. Conley, Lawrence Berkeley National Laboratory,

report that an optical spectrum (range 330–900 nm) of SN 2000ca (IAUC 7409), taken with the Cerro Tololo 4-m telescope (+ RC spectrograph) on Apr. 29.3 UT, exhibits the features of a type-Ia supernova near maximum light. The spectrum shows a strong Si II 635.5-nm feature; the photospheric expansion velocity is 11 000 km/s, based on the minimum of this feature and the NED redshift for the host galaxy ESO 383-32 (z = 0.024). The ratio of Si II features at 580 and 610 nm is ∼ 0.18, suggesting that this supernova is slightly hotter and more luminous than normal (cf. Nugent et al. 1995, Ap.J. 455, L147). Very weak absorption lines (EW ∼ 0.02 nm) due to Na D at the redshift of the host galaxy may be present, indicating modest extinction by the host galaxy.

SUPERNOVA 2000E IN NGC 6951 CCD R magnitudes: Apr. 21.007 UT, 16.1 ± 0.1 (K. Hornoch, Leleko-

vice, Czech Republic); 21.85, 16.0 ± 0.1 (Hornoch); 22.862, 16.1 ± 0.1 (Hornoch); 23.866, 16.0 ± 0.1 (Hornoch); 24.074, 16.0 ± 0.1 (D. Hanžl, Brno, Czech Rep.); 24.838, 15.9 ± 0.2 (Hornoch); 28.991, 16.0 ± 0.1 (Hornoch); May 1.087, 16.2 ± 0.1 (Hanžl).

SUPERNOVA 2000cb IN IC 1158

Corrigendum. On IAUC 7410, line 25, for 10 000 km/s. read 18 200 km/s.

2000 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green